# Bound states for systems with quartic energy-momentum dispersion

### v1

@@ V Finite square well potential
$\psi_{II}(x)=C_{1}\cos(px)+C_{2}\cosh(px),\qquad p=\frac{E^{1/4}}{\hbar a},\qquad 0<x<L.$ (32)
@@
for states of positive and negative parity, respectively. In contrast to the Schrödinger equation in the infinite well limit, Eq.(50) is transcendental and corresponds to the infinite well problem with the boundary conditions $\psi(\pm L)=0$ and $\psi^{\prime}(\pm L)=0$. As in the case of Schrödinger’s equation, the energy levels in an infinite well potential are higher than the corresponding values for a finite well potential because the wave function in the latter case penetrates into a classically forbidden region. Further, the wave function of the ground state is nodeless. In fact, for systems with quartic dispersion, it was proved [35, 36] that the node theorem is valid for infinite well $v\to\infty$. However, the wave function for a finite potential well problem has infinite number of oscillations in the classically forbidden region confirming our conclusion made in Sec.IV. It is worth mentioning also that, for a finite width barrier, oscillations of wave functions may lead to the appearance of oscillating tunneling current [37, 38].
@@
Equations (38) and (48) (or, equivalently, Eqs.(40) and 49) are eigenvalue equations for $pL=\varepsilon^{1/4}L$ defining it as a function of a dimensionless combination $vL^{4}$ of the depth $v$ and width $L$ of the potential well because $kL=(vL^{4}-(pL)^{4})^{1/4}$. The behaviour of $pL$ as a function of $\lambda=v^{1/4}L$ for states of both parities is shown in Fig.4. It is not difficult to show that the number of eigenvalues $N(v,L)$ of both parities grows asymptotically as $\simeq\lambda/\pi$ for $\lambda\to\infty$. It is clearly seen that all levels, except the ground state, start with some threshold value $\lambda_{th}$. The ground state is of positive parity and exists for arbitrary shallow potential well, i.e., for any $v>0$ like in the case of the one-dimensional Schrödinger’s equation.
@@
VI Summary
@@
To investigate bound states for quasiparticles in weakly dispersing energy bands, we studied 1D systems with the quartic energy-momentum dispersion and the quartic potential (the double quartic problem). To determine bound state energies, we applied the WKB semiclassical method formulated for systems with quartic dispersion in [20]. Since the connection formulas in the vicinity of turning points require the use of hyperasymptotics of higher-order Airy functions for fourth order differential equations, the Bohr-Sommerfeld quantization rule for systems with quartic dispersion contains non-perturbative in the Planck constant contribution.
To analyse the role of higher order WKB corrections, we calculated the second and fourth WKB corrections and found that like the non-perturbative contribution these higher order WKB corrections are mostly relevant for the lowest bound states energies with their contributions quickly decreasing for states with higher energy. To quantify the accuracy of the energy levels found using the WKB approach, we computed the bound state energy numerically by using the variational method with the universal Gaussian basis. Quantitatively, we found that the WKB corrections notably improve the accuracy of WKB results and the deviation of the energy obtained in the WKB approach from the numerically calculated energy is around 8 percent for the lowest energy state. For states with higher energy this deviation quickly decreases.
@@
As to wave functions of bound states, they provide important information also. Our main principal finding is the result that the wave function of the bound states have nodes (in fact, infinitely many) in the classically forbidden region for systems with quartic dispersion unlike the Schrödinger equation where the bound state wave functions do not have such nodes. On the other hand, the number of nodes in the classically allowed region coincides with the quantum number $n$ of the corresponding bound state. To check this observation, we considered an exactly solvable potential well problem and found that the wave function of the ground state indeed does not have nodes in the classically allowed region and oscillates producing nodes in regions beyond the potential well. The higher energy states also have nodes in the classically forbidden region.
@@
Interestingly, although our findings show that the oscillation theorem is not valid for systems with quartic dispersion, this theorem holds in the classically allowed region where the number of nodes of the wave function coincides with the quantum number $n$ of the corresponding bound state. The mathematical reason for the presence of nodes of the ground state wave function in the classically forbidden region is the existence of both real and imaginary parts in all four semiclassical momenta defining the asymptotics of the wave function in the classically forbidden region. This is in contrast to the case of Schrödinger‘s equation where the semiclassical momentum is purely imaginary in the classically forbidden region which gives non-oscillating exponentially decreasing asymptotics. Oscillations in the classically forbidden region may have important physical consequences, for example, lead to oscillating tunneling current [37, 38].

### v2

@@ IV Spectrum and wave function nodes for square well potential
$\psi_{II}(x)=C_{1}\cos(px)+C_{2}\cosh(px),\qquad p=\frac{E^{1/4}}{\hbar a}.$ (27)
@@
for states of positive and negative parity, respectively. In contrast to the Schrödinger equation in the infinite well limit, Eq.(45) is transcendental and corresponds to the infinite well problem with the boundary conditions $\psi(\pm L)=0$ and $\psi^{\prime}(\pm L)=0$. As in the case of Schrödinger’s equation, the energy levels in an infinite well potential are higher than the corresponding values for a finite well potential because the wave function in the latter case penetrates into a classically forbidden region. Further, the wave function of the ground state is nodeless. In fact, for systems with quartic dispersion, it was proved [26, 27] that the node theorem is valid for infinite well $v\to\infty$. However, the wave function for a finite potential well problem has infinite number of oscillations in the classically forbidden region confirming our conclusion made in Sec.III. It is worth mentioning also that, for a finite width barrier, oscillations of wave functions may lead to the appearance of oscillating tunneling current [28, 29].
@@
Equations (33) and (43) (or, equivalently, Eqs.(35) and 44) are eigenvalue equations for $pL=\varepsilon^{1/4}L$ defining it as a function of a dimensionless combination $vL^{4}$ of the depth $v$ and width $L$ of the potential well because $kL=(vL^{4}-(pL)^{4})^{1/4}$. The behaviour of $pL$ as a function of $\lambda=v^{1/4}L$ for states of both parities is shown in Fig.4. It is not difficult to show that the number of eigenvalues $N(v,L)$ of both parities grows asymptotically as $\simeq\lambda/\pi$ for $\lambda\to\infty$. It is clearly seen that all levels, except the ground state, start with some threshold value $\lambda_{th}$. The ground state is of positive parity and exists for arbitrary shallow potential well, i.e., for any $v>0$ like in the case of the one-dimensional Schrödinger’s equation.
@@
Thus, the analysis of eigenfunctions of an exactly solvable square well potential problem considered in this section confirms our proposition that, for one-dimensional Hamiltonians with quartic dispersion, the node-counting rule holds only in the classically allowed region.
V Summary
@@
In the present paper, we studied the properties of bound-state wave functions and bound-state energies of quasiparticles in weakly dispersing energy bands for 1D systems with the quartic energy-momentum dispersion and the quartic potential (the double quartic problem). To determine bound state energies, we applied the WKB semiclassical method formulated for systems with quartic dispersion in [8] which gives the Bohr-Sommerfeld quantization rule with non-perturbative in the Planck constant contribution. We calculated the second and fourth WKB corrections and shown that the non-perturbative contribution, as well as these higher order WKB corrections are mostly relevant for the lowest bound states energies with their contributions quickly decreasing for states with higher energy.
@@
To check the accuracy of our results found using the WKB approach, we computed the bound state energy numerically by using the variational method with the universal Gaussian basis. Our main principal finding is that the classical oscillation theorem fails in the classically forbidden region for systems with quartic energy-momentum dispersion. We found that the bound-state wave functions have nodes (in fact, infinitely many) in the classically forbidden region for systems with quartic dispersion unlike the Schrödinger equation where the bound state wave functions do not have such nodes. On the other hand, the number of nodes in the classically allowed region coincides with the quantum number $n$ of the corresponding bound state. To check this observation, we considered an exactly solvable potential well problem and found that the wave function of the ground state indeed does not have nodes in the classically allowed region and oscillates producing nodes in regions beyond the potential well. The higher energy states also have nodes in the classically forbidden region whose number equals $n$. Hence, the oscillation theorem holds in the classically allowed region in all considered examples.
@@
The mathematical reason for the presence of nodes of the ground state wave function in the classically forbidden region is the existence of both real and imaginary parts in all four semiclassical momenta defining the asymptotics of the wave function in the classically forbidden region. This is in contrast to the case of Schrödinger‘s equation where the semiclassical momentum is purely imaginary in the classically forbidden region which gives non-oscillating exponentially decreasing asymptotics. Oscillations in the classically forbidden region may have important physical consequences, for example, lead to oscillating tunneling current [28, 29].